OpenStudy (anonymous):

find an equation of the line having the slope and containg the given point m=3/4,(3,-2)

OpenStudy (anonymous):

hes apparently in a exam @lgbasallote

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

we can write (y-(-2))=(3/4)(x-3) y=3x/4-1/4

OpenStudy (lgbasallote):

let's give him the benefit of the doubt...just tell him HOW to solve it..

OpenStudy (anonymous):

thanks drewnj

OpenStudy (anonymous):

find an equation of the line containing the points (4,1) and (12,8)

OpenStudy (anonymous):

How about you use the method shown above^^ by @AndrewNJ ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Just find your slope and then use the method shown.

OpenStudy (anonymous):

first u have to find out the slpoe k k=(y1-y2)/(x1-x2)=(8-1)/(12-4)=7/8 then substitution the slope k and one point [(4,1) or (12,8) ]in to a basic quadratic equation y=kx+b if i use (4,1) then 1=(7/8)*4+b we can work out b=-5/2 so y=(7/8)x-5/2

OpenStudy (anonymous):

u are right! the work you done can have the slope can substitution it and a point into y=kx+b to have the b

OpenStudy (anonymous):

the second 'can' should be 'and' -.-

OpenStudy (anonymous):

so -7=(5/7)*0+b right

OpenStudy (anonymous):

you reverse x and y it should be 0=(5/7)*(-7)+b remember y=kx+b

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so it should read y=(5/7)x-1

OpenStudy (anonymous):

0=(5/7)*(-7)+b 0=-5+b b=5

OpenStudy (anonymous):

so it should read y=(5/7)x+5 right

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok i am getting it

OpenStudy (anonymous):

what's your mean for it ?

OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line

OpenStudy (anonymous):

rewrite 2x+9y=4 to y=(-2/9)x+4/9 and (-2/9) is the slope

OpenStudy (anonymous):

no it just work out the slope u still need to sub the slope and the point (0,7) into y=kx+b to find b

OpenStudy (anonymous):

couldnt you also write it as 2(0)+9(7)=4

OpenStudy (anonymous):

oh im sorry the upper one is perpendicular, i thought it was Parallel i redo it again

OpenStudy (anonymous):

rewrite 2x+9y=4 to y=(-2/9)x+4/9 and the slope of that line is -2/9 so the slope perpendicular to that line is k=1/(-2/9)=-9/2 and then substitute k=-9/2 and point (0,7) into y=kx+b to find b

OpenStudy (anonymous):

so the equation should read 0=-9/2*7+b

OpenStudy (anonymous):

this time the point is (0,7) x=0, y=7. so 7=(-9/2)*0+b

OpenStudy (anonymous):

oh yea thats right

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so next line should read 7=0+b then y=-7+b

OpenStudy (anonymous):

so we work out b=7 and FINALLY substitute k=-9/2 and b=7 into y=kx+b the final equation is y=(-9/2)x+7

OpenStudy (anonymous):

ok whatever you say

OpenStudy (anonymous):

you a smart dude

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what grade you in